OpenStudy (anonymous):

Prove this:

OpenStudy (anonymous):

\[\large \sin(x+y)\sin(x-y)=\sin^2x-\sin^2y\]

hartnn (hartnn):

ok, do u know the formula of 2 sinA sin B =...?

OpenStudy (anonymous):

\[ LS=(sinxsosx+sinycosy)(sixcosx-sinycosy)\]

hartnn (hartnn):

i think u need to revise your formulas....how'd u get that ?

hartnn (hartnn):

sin(x+y) = sin x cosy + cos x sin y

OpenStudy (anonymous):

sin (x+y) formula and then the sin (x-y)

hartnn (hartnn):

sin(x-y) = sin x cosy - cos x sin y

hartnn (hartnn):

but here its better to use 2sin A sin B =... ? formula....

OpenStudy (anonymous):

no wonder -.-

OpenStudy (anonymous):

2sinxcosx

hartnn (hartnn):

\(\large 2sinAsinB=cos(A-B)-cos(A+B)\) use this. A=x+y, B=x-y and tell me what u get ?

OpenStudy (anonymous):

ohh we didnt use that one

OpenStudy (anonymous):

is this for LS ?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

not sure but, = (cosxcosy+sinxsiny) - ( cosxcosy-sinxsiny )

hartnn (hartnn):

i think you are not comfortable with that formula, lets use the formula you know sin(x+y) = sin x cosy + cos x sin y sin(x-y) = sin x cosy - cos x sin y so multiplying the above 2 equations , you get LS = sin (x+y) sin (x-y) =[ sin x cosy + cos x sin y ] [sin x cosy - cos x sin y] =..... can you figure out next step ??

hartnn (hartnn):

its of the form [a+b][a-b] =....?

OpenStudy (anonymous):

\[=\sin^2xcos^2y-sinxsinycosxcosy+cosxcosysinxsiny-\cos^2xsin^2y\]

hartnn (hartnn):

yeah, that simplifies to ... ?

hartnn (hartnn):

does any term gets cancelled ?

OpenStudy (anonymous):

the middle two

hartnn (hartnn):

correct! now in RS you only want sine terms, right ? so convert cos^2 y and cos^2 x into sine terms, do u know how ??

OpenStudy (anonymous):

wait on the left side = \[\sin^2xcos^2x-\cos^2xsin^2y\]

hartnn (hartnn):

its \(\sin^2xcos^2y-\cos^2xsin^2y\)

OpenStudy (anonymous):

opps typo okay :)

hartnn (hartnn):

so now convert all cos to sin...

OpenStudy (anonymous):

\[\sin^2x(1-\sin^2y)-(1-\sin^2x)\sin^2y\]

hartnn (hartnn):

absolutely correct if you simplify , you directly get RS try it.

OpenStudy (anonymous):

Yup, i got it ! thankyou you are a great teacher ! :)

hartnn (hartnn):

welcome ^_^